# PointRecon: Online Point-based 3D Reconstruction via Ray-based 2D-3D Matching

### v1

@@ Abstract
We propose a novel online, point-based 3D reconstruction method from posed monocular RGB videos. Our model maintains a global point cloud representation of the scene, continuously updating the features and 3D locations of points as new images are observed. It expands the point cloud with newly detected points while carefully removing redundancies. The point cloud updates and the depth predictions for new points are achieved through a novel ray-based 2D-3D feature matching technique, which is robust against errors in previous point position predictions. In contrast to offline methods, our approach processes infinite-length sequences and provides real-time updates. Additionally, the point cloud imposes no pre-defined resolution or scene size constraints, and its unified global representation ensures view consistency across perspectives. Experiments on the ScanNet dataset show that our method achieves state-of-the-art quality among online MVS approaches. Project page: https://arthurhero.github.io/projects/pointrecon/
@@ 1 Introduction
Deep learning has enabled a wide range of MVS methods. Depth prediction-based approaches (Kendall et al., 2017; Chang & Chen, 2018; Yao et al., 2018; Wang & Shen, 2018; Huang et al., 2018; Duzceker et al., 2021; Im et al., 2019; Sayed et al., 2022) accumulate features from multiple views into a per-view cost volume to regress depth maps, while volumetric methods (Kar et al., 2017; Murez et al., 2020; Sun et al., 2021; Stier et al., 2021; Bozic et al., 2021; Gao et al., 2023) aggregate features into a global voxel grid to predict occupancy or TSDF values. Point cloud-based methods (Lhuillier & Quan, 2005; Chen et al., 2019) directly estimate surface point positions. However, depth prediction methods regress depth maps independently for each image, which can lead to inconsistencies across views. Volumetric methods maintain a global voxel grid, but require predefined voxel sizes and suffer from high memory costs at high resolutions, resulting in coarse meshes. Additionally, they are limited to bounded areas and cannot extend to an infinite horizon, despite the advantages of using cameras.
A globally consistent 3D point cloud representation addresses some of these drawbacks. It is a sparse representation that significantly reduces memory consumption compared to volumetric methods. Furthermore, point clouds do not require predefined voxel sizes and can adaptively represent surface details with varying densities, allowing for greater detail in areas that need it. However, existing point cloud-based MVS methods typically rely on iterative optimization (Chen et al., 2019; Kerbl et al., 2023), making them unsuitable for online algorithms.
@@
It is important to note that the scope of online MVS methods differs from that of SLAM methods (Mur-Artal et al., 2015; Teed & Deng, 2021; Zhang et al., 2023; Zhu et al., 2024; Zhou et al., 2024). Our method recovers geometry from a sequence of posed images and operates purely in a feed-forward manner, while SLAM methods estimate both camera pose and geometry, typically involving optimization steps like the Gauss-Newton algorithm and bundle adjustment.
@@
  • We propose an online point-based MVS method that maintains a feature-augmented 3D point cloud, unconstrained by predefined voxel resolutions, scene sizes or input sequence length.
@@
  • Our results on the popular ScanNetv2 dataset demonstrate that our approach achieves state-of-the-art quality among online methods while providing greater detail than volumetric approaches.
@@ 2 Related Work
Depth Prediction via Multi-view Stereo Matching. Depth can be inferred from image patch similarity between two posed cameras viewing the same object (Newcombe et al., 2011; Pizzoli et al., 2014), but manually designed measures are often unreliable. CNN-based methods (Zbontar & LeCun, 2016; Luo et al., 2016) improve this by learning patch similarity, though they still lack global context and require post-processing. GC-Net (Kendall et al., 2017) introduces cost volumes and 3D convolutions for global context, while PSM-Net (Chang & Chen, 2018) further improves it with multi-scale feature maps. DeepMVS (Huang et al., 2018) and MVSNet (Yao et al., 2018), both extend stereo matching to multiple views, with DeepMVS using plane-sweep volumes and MVSNet introducing a camera frustum-based 3D cost volume. MVDepthNet (Wang & Shen, 2018) further refines this process by compressing features into a single volume for efficient 2D convolution. DPSNet (Im et al., 2019) improves results by matching deeper features, DeepVideoMVS (Duzceker et al., 2021) integrates reconstruction history with a recurrent network, and SimpleRecon (Sayed et al., 2022) further enhances depth map quality by incorporating geometric metadata into the cost volume. While depth map-based MVS methods break down large scenes for faster inference, without a unified scene representation, the independently predicted depth maps fail to ensure consistent views.
@@
Volumetric TSDF Regression. Instead of predicting depth maps, some methods directly generate global surfaces from cost volumes. LSM (Kar et al., 2017) back-projects features into a global voxel grid for each image and fuses them to regress global voxel occupancy. Atlas (Murez et al., 2020) extends this to scenes by accumulating features into a global voxel grid and using 3D convolution to regress TSDF values. NeuralRecon (Sun et al., 2021) introduces an online approach that incrementally constructs and fuses local grids with a global grid using a GRU, while also sparsifying fine grids based on coarse predictions. Building on this, TransformerFusion (Bozic et al., 2021) leverages transformers to selectively fuse image features into the voxel grid. Further improvements are made by VisFusion (Gao et al., 2023), which incorporates ray-based sparsification, and DG-Recon (Ju et al., 2023), which enhances the process with monocular depth priors. More recent offline methods like VoRTX (Stier et al., 2021) incorporate ray direction and depth for view-aware attention, while FineRecon (Stier et al., 2023) and CVRecon (Feng et al., 2023) enhance results through fine-grained supervision and cost-volume integration. While volumetric methods offer consistency and can infer unseen surfaces, they rely on predefined grids (often 4cm in size), limiting their ability to capture finer details.
@@
Point Cloud-based Reconstruction. Early methods like (Lhuillier & Quan, 2005) extract keypoints, back-project them into 3D space to form a point cloud, and optimize a surface over it. Deep learning-based approaches like Point-MVSNet (Chen et al., 2019) predict a coarse depth map, back-project to a point cloud, augment it with multi-view features, and iteratively refine point positions or depth maps. However, these methods are not online and focus primarily on object datasets. To the best of our knowledge, our method is the first online point-based MVS approach generalized to large scenes. Recently, Gaussian splatting-based reconstruction methods have gained attention, where each Gaussian can be viewed as a point with scale and opacity. While the original 3D GS (Kerbl et al., 2023) requires per-scene optimization, recent methods have extended it to feed-forward reconstruction Charatan et al. (2024); Tang et al. (2024); Chen et al. (2024); Zhang et al. (2024). In particular, Long-LRM (Ziwen et al., 2024) achieves scene-level GS reconstruction from long-sequence input in a fully feed-forward manner. However, these methods mainly focus on color rendering from novel views, whereas the primary goal of MVS is to reconstruct geometry. Without explicit depth supervision, Gaussian-based methods often suffer from the “floater” issue, where translucent Gaussian clouds appear at incorrect locations. Our method could serve as a strong initialization for future online, feed-forward, large-scene Gaussian reconstructions.
@@ 3.1 Backbone, Feature Pyramid and Monocular Depth Prediction
We first pass the image $\mathbf{I}_{t}$ through an image encoder to produce a set of multi-level feature maps $\{\mathbf{F}^{l}\}_{l=1}^{4}$, where level 1 is the coarsest and level 4 the finest. While our approach is compatible with any image encoder, such as ResNets (He et al., 2016), we specifically use AutoFocusFormer (Ziwen et al., 2023), a transformer-based encoder that excels at preserving key details by automatically locating important points during downsampling. This results in non-uniform 2D point cloud feature maps, represented as $\mathbf{F}^{l}=\{(p^{\circ}_{1},f_{1}),(p^{\circ}_{2},f_{2}),\dots\}$ at each level.
@@ 3.2 Scene Update
$p_{1}(r_{i},r_{j}^{\circ})=\arg\min_{p\in\mathcal{R}_{i}}\min_{p^{\prime}\in\mathcal{R}_{j}^{\circ}}\|p-p^{\prime}\|$ (1)
@@ 3.3 Depth Prediction
We again compute the dot product between the reduced 2D and 3D features. As in the scene update step, we append geometric metadata to the dot products. The core idea is that if the 2D point $p^{\circ}_{i}$ matches with a 3D ray $r_{j}$, then the depth of the crossing between $r_{i}^{\circ}$ and $r_{j}$ is likely the correct depth for $p^{\circ}_{i}$. Concretely, for a 2D point $p_{i}^{\circ}$ with camera ray $r_{i}^{\circ}$, a sampled position $s_{j}$ on $r_{i}^{\circ}$, and a 3D neighbor with position $p_{j}$ and ray $r_{j}$, the list of geometric metadata includes:
@@
Similar to the scene update step, we concatenate the metadata with the reduced 3D features and the dot products at each sampled position, passing them through a small MLP. The outputs from all sampled positions are then combined with the reduced feature of $p^{\circ}_{i}$, and this combined vector is processed by another small MLP to produce $h_{i}$, which represents the information $p^{\circ}_{i}$ has gained from the scene point cloud.
@@ 3.4 Point Cloud Merging
$d_{i}=\sum_{P(\mathbf{q}_{j},I_{t})=p_{i}^{\circ}}\text{softmax}(c_{j})d_{j}$ (3)
@@
$d_{i}=\min_{\begin{subarray}{c}P(\mathbf{q}_{j},\mathbf{I}_{t})=p_{i}^{\circ}\\ c_{j}>\epsilon\end{subarray}}d_{j}$ (4)
@@
Note that while we train the model using 9-view local windows, during evaluation, our goal is to seamlessly reconstruct the entire scene. As the sequence length increases, the number of cameras sharing a partial view can reach into the hundreds, leading to a significant computational burden. To address this, we employ a heuristic: if a camera is not among the $K$ most recent cameras in the current merging step, we retain all points originating from that camera, preventing their deletion, and remove the camera from the current merging algorithm. We empirically set $K=16$.
@@ 3.5 Loss Functions
We train our model by supervising the depth maps produced by different modules of our model: the monocular depth prediction, the depth prediction with feature matching and the depth maps rendered at the point merging step. The three types of depth maps use the same set of losses: an $L1$ depth loss, a gradient loss and a normal loss following  (Sayed et al., 2022).
@@
$\mathcal{L}_{\text{depth}}=\dfrac{1}{HW}\sum_{l=1}^{4}\dfrac{1}{l^{2}}\sum_{i}|\log d_{i}^{l}-\log d_{i}^{\text{gt}}|$ (5)
@@
$\mathcal{L}_{\text{grad}}=\dfrac{1}{HW}\sum_{r=0}^{3}\sum_{i}|\nabla d_{i\downarrow_{2^{r}\times}}^{4}-\nabla d_{i\downarrow_{2^{r}\times}}^{\text{gt}}|$ (6)
@@
$\mathcal{L}_{\text{normal}}=\dfrac{1}{2HW}\sum_{i}1-\mathbf{N}_{i}\cdot\mathbf{N}_{i}^{\text{gt}}$ (7)
@@
$\mathcal{L}_{\text{update}}=\dfrac{1}{N}\sum_{l=1}^{4}|z_{i}-z_{i}^{\text{gt}}|$ (8)
@@
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{depth}}+\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{normal}}+\mathcal{L}_{\text{update}}$ (9)
@@ 4.1 Implementation Details
We select AFF-Mini, the smallest variant from Ziwen et al. (2023), which consists of only 6.75M parameters, as our image encoder backbone. Our model is trained and evaluated using the ScanNetv2 (Dai et al., 2017) dataset. For input images, we resize the resolution to $640\times 480$. The depth maps are predicted at a resolution of $160\times 120$.
@@
Training details. We begin with training the model on image pairs for 2 epochs, using the two images to mutually serve as sources for feature matching, without any point merging at this stage. Then, we train the model on 9-view local image windows for 6 epochs, feeding the images sequentially into the model and performing point merging at each time step. We use a learning rate $10^{-4}$ for the first $4$ epochs, reducing it by a factor of 0.1 at the fifth epoch and the eighth epoch. We train with the AdamW optimizer (Loshchilov & Hutter, 2017) with a weight decay of $10^{-4}$.
During training, we apply random color augmentation to the input RGB images with a probability of 0.8, using TorchVision (Marcel & Rodriguez, 2010). The color augmentation parameters include brightness set to 0.4, contrast to 0.4, saturation to 0.2, and hue to 0.1. Additionally, we randomly reverse the order of the input sequences with a probability of 0.5. For keyframe selection, we follow the methodology outlined in (Duzceker et al., 2021). We reduce the weight for the loss of monocular depth prediction by a factor of 0.5 at the fifth and eighth epochs.
Evaluation details. During evaluation, we sequentially feed the keyframe images into the model, performing point merging at each time step. Once the point cloud for the entire scene is constructed, we render depth maps using the keyframe cameras by projecting the points onto the image planes. Each pixel in the rendered depth maps corresponds to the depth of the closest point (as detailed in Sec. 3.4). To construct a 3D mesh from the rendered depth maps, we apply TSDF fusion (Curless & Levoy, 1996).

### v2

@@ Abstract
We propose a novel online, point-based 3D reconstruction method from posed monocular RGB videos. Our model maintains a global point cloud representation of the scene, continuously updating the features and 3D locations of points as new images are observed. It expands the point cloud with newly detected points while carefully removing redundancies. The point cloud updates and the depth predictions for new points are achieved through a novel ray-based 2D-3D feature matching technique, which is robust against errors in previous point position predictions. In contrast to offline methods, our approach processes infinite-length sequences and provides real-time updates. Additionally, the point cloud imposes no pre-defined resolution or scene size constraints, and its unified global representation ensures view consistency across perspectives. Experiments on the ScanNet dataset show that our method achieves comparable quality among online MVS approaches. Project page: https://arthurhero.github.io/projects/pointrecon/
@@ 1 Introduction
Deep learning has enabled a wide range of MVS methods. Depth prediction-based approaches [16, 2, 34, 33, 12, 8, 13, 26] accumulate features from multiple views into a per-view cost volume to regress depth maps, while volumetric methods [15, 23, 30, 28, 1, 10] aggregate features into a global voxel grid to predict occupancy or TSDF values. Point cloud-based methods [18, 4] directly estimate surface point positions. However, depth prediction methods regress depth maps independently for each image, which can lead to inconsistencies across views. Volumetric methods maintain a global voxel grid, but require predefined voxel sizes and suffer from high memory costs at high resolutions, resulting in coarse meshes. Additionally, they are limited to bounded areas and cannot extend to an infinite horizon, despite the advantages of using cameras.
A globally consistent 3D point cloud representation addresses some of these drawbacks. It is a sparse representation that significantly reduces memory consumption compared to volumetric methods. Furthermore, point clouds do not require predefined voxel sizes and can adaptively represent surface details with varying densities, allowing for greater detail in areas that need it. However, existing point cloud-based MVS methods typically rely on iterative optimization [4, 17], making them unsuitable for online algorithms.
@@
It is important to note that the scope of online MVS methods differs from that of SLAM methods [22, 32, 37, 39, 38]. Our method recovers geometry from a sequence of posed images and operates purely in a feed-forward manner, while SLAM methods estimate both camera pose and geometry, typically involving optimization steps like the Gauss-Newton algorithm and bundle adjustment.
@@
  • We propose an online point-based MVS method that maintains a feature-augmented 3D point cloud, unconstrained by input sequence length, predefined voxel resolutions or scene sizes, while ensuring consistency among viewpoints.
@@
  • Our results on the popular ScanNetv2 dataset demonstrate that our approach achieves comparable quality to existing online methods while providing the flexibility mentioned above.
@@ 2 Related Work
Depth Prediction via Multi-view Stereo Matching. Depth can be inferred from image patch similarity between two posed cameras viewing the same object [24, 25], but manually designed measures are often unreliable. CNN-based methods [35, 20] improve this by learning patch similarity, though they still lack global context and require post-processing. GC-Net [16] introduces cost volumes and 3D convolutions for global context, while PSM-Net [2] further improves it with multi-scale feature maps. DeepMVS [12] and MVSNet [34], both extend stereo matching to multiple views, with DeepMVS using plane-sweep volumes and MVSNet introducing a camera frustum-based 3D cost volume. MVDepthNet [33] further refines this process by compressing features into a single volume for efficient 2D convolution. DPSNet [13] improves results by matching deeper features, DeepVideoMVS [8] integrates reconstruction history with a recurrent network, and SimpleRecon [26] further enhances depth map quality by incorporating geometric metadata into the cost volume. While depth map-based MVS methods break down large scenes for faster inference, without a unified scene representation, the independently predicted depth maps fail to ensure consistent views.
@@
Volumetric TSDF Regression. Instead of predicting depth maps, some methods directly generate global surfaces from cost volumes. LSM [15] back-projects features into a global voxel grid for each image and fuses them to regress global voxel occupancy. Atlas [23] extends this to scenes by accumulating features into a global voxel grid and using 3D convolution to regress TSDF values. NeuralRecon [30] introduces an online approach that incrementally constructs and fuses local grids with a global grid using a GRU, while also sparsifying fine grids based on coarse predictions. Building on this, TransformerFusion [1] leverages transformers to selectively fuse image features into the voxel grid. Further improvements are made by VisFusion [10], which incorporates ray-based sparsification, and DG-Recon [14], which enhances the process with monocular depth priors. More recent offline methods like VoRTX [28] incorporate ray direction and depth for view-aware attention, while FineRecon [29] and CVRecon [9] enhance results through fine-grained supervision and cost-volume integration. While volumetric methods offer consistency and can infer unseen surfaces, they rely on predefined grids (often 4cm in size), limiting their ability to capture finer details.
@@
Point Cloud-based Reconstruction. Early methods like [18] extract keypoints, back-project them into 3D space to form a point cloud, and optimize a surface over it. Deep learning-based approaches like Point-MVSNet [4] predict a coarse depth map, back-project to a point cloud, augment it with multi-view features, and iteratively refine point positions or depth maps. However, these methods are not online and focus primarily on object datasets. To the best of our knowledge, our method is the first online point-based MVS approach generalized to large scenes. Recently, Gaussian splatting-based reconstruction methods have gained attention, where each Gaussian can be viewed as a point with scale and opacity. While the original 3D GS [17] requires per-scene optimization, recent methods have extended it to feed-forward reconstruction [3, 31, 5, 36]. In particular, Long-LRM [41] achieves scene-level GS reconstruction from long-sequence input in a fully feed-forward manner. However, these methods mainly focus on color rendering from novel views, whereas the primary goal of MVS is to reconstruct geometry. Without explicit depth supervision, Gaussian-based methods often suffer from the “floater” issue, where translucent Gaussian clouds appear at incorrect locations. Our method could serve as a strong initialization for future online, feed-forward, large-scene Gaussian reconstructions.
@@ 3.1 Backbone, Feature Pyramid and Monocular Depth Prediction
We first pass the image $\mathbf{I}_{t}$ through an image encoder to produce a set of multi-level feature maps $\{\mathbf{F}^{l}\}_{l=1}^{4}$, where level 1 is the coarsest and level 4 the finest. While our approach is compatible with any image encoder, such as ResNets [11], we specifically use AutoFocusFormer [40], a transformer-based encoder that excels at preserving key details by automatically locating important points during downsampling. This results in non-uniform 2D point cloud feature maps, represented as $\mathbf{F}^{l}=\{(p^{\circ}_{1},f_{1}),(p^{\circ}_{2},f_{2}),\dots\}$ at each level.
@@ 3.2 Scene Update
$\small p_{1}(r_{i},r_{j}^{\circ})=\arg\min_{p\in\mathcal{R}_{i}}\min_{p^{\prime}\in\mathcal{R}_{j}^{\circ}}\|p-p^{\prime}\|$ (1)
@@ 3.3 Depth Prediction
We again compute the feature dot product between the reduced 2D and 3D features. As in the scene update step, we append geometric metadata to the dot products. The core idea is that if the 2D point $p^{\circ}_{i}$ matches with a 3D ray $r_{j}$, then the depth of the crossing between $r_{i}^{\circ}$ and $r_{j}$ is likely the correct depth for $p^{\circ}_{i}$. Concretely, for a 2D point $p_{i}^{\circ}$ with camera ray $r_{i}^{\circ}$, a sampled position $s_{j}$ on $r_{i}^{\circ}$, and a 3D neighbor with position $p_{j}$ and ray $r_{j}$, the list of geometric metadata includes:
@@
Similar to the scene update step, we concatenate the metadata with the reduced 3D features and the feature dot products at each sampled position, passing them through a small MLP. The outputs from all sampled positions are then combined with the reduced feature of $p^{\circ}_{i}$, and this combined vector is processed by another small MLP to produce $h_{i}$, which represents the information $p^{\circ}_{i}$ has gained from the scene point cloud.
@@ 3.4 Point Cloud Merging
$\small d_{i}=\sum_{P(\mathbf{q}_{j},I_{t})=p_{i}^{\circ}}\text{softmax}(c_{j})d_{j}$ (3)
@@
$\small d_{i}=\min_{\begin{subarray}{c}P(\mathbf{q}_{j},\mathbf{I}_{t})=p_{i}^{\circ}\\ c_{j}>\epsilon\end{subarray}}d_{j}$ (4)
@@
Note that while we train the model using 9-view local windows, our goal is to seamlessly reconstruct the entire scene during evaluation. As the sequence length increases, the number of cameras sharing a partial view can reach into the hundreds, leading to a significant computational burden. To address this, we employ a heuristic: if a camera is not among the $K$ most recent cameras in the current merging step, we retain all points covering the pixels of that camera, preventing their deletion, and remove the camera from the current merging algorithm. We empirically set $K=16$.
@@ 3.5 Loss Functions
We train our model by supervising the depth maps produced by different modules of our model: the monocular depth prediction, the depth prediction with feature matching and the depth maps rendered at the point merging step. The three types of depth maps use the same set of losses: an $L1$ depth loss, a gradient loss and a normal loss following  [26].
@@
$\small\mathcal{L}_{\text{depth}}=\dfrac{1}{HW}\sum_{l=1}^{4}\dfrac{1}{l^{2}}\sum_{i}|\log d_{i}^{l}-\log d_{i}^{\text{gt}}|$ (5)
@@
$\small\mathcal{L}_{\text{grad}}=\dfrac{1}{HW}\sum_{r=0}^{3}\sum_{i}|\nabla d_{i\downarrow_{2^{r}\times}}^{4}-\nabla d_{i\downarrow_{2^{r}\times}}^{\text{gt}}|$ (6)
@@
$\small\mathcal{L}_{\text{normal}}=\dfrac{1}{2HW}\sum_{i}1-\mathbf{N}_{i}\cdot\mathbf{N}_{i}^{\text{gt}}$ (7)
@@
$\small\mathcal{L}_{\text{update}}=\dfrac{1}{N}\sum_{l=1}^{4}|z_{i}-z_{i}^{\text{gt}}|$ (8)
@@
$\small\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{depth}}+\mathcal{L}_{\text{grad}}+\mathcal{L}_{\text{normal}}+\mathcal{L}_{\text{update}}$ (9)
[…]
Offline COLMAP [27] / 0.264 0.137 0.138 - 83.4 87.1
Atlas [23] / 0.123 0.065 0.045 - 93.6 99.9
VoRTX [28] / 0.092 0.058 0.036 - 93.8 95.0
CVRecon [9] / 0.078 0.047 0.028 - 96.3 -
FineRecon [29] / 0.069 0.042 0.026 86.6 97.1 97.2
Online NeuralRecon [30] 90 0.106 0.065 0.031 - 94.8 90.9
TF [1] 326 0.099 0.065 0.042 - 93.4 90.5
SimpleRecon [26] 72 0.083 0.046 0.022 - 95.4 94.4
PointRecon (4cm) 618 0.085 0.054 0.022 71.9 96.4 94.8
PointRecon (2cm) 618 0.087 0.055 0.024 72.1 96.2 94.6
Table 1: Depth Map Evaluation. Depth map quality for the ScanNetv2 test split, rendered from the reconstructed 3D meshes. We follow the evaluation protocol of Atlas [23]. - indicates that the metric was not reported in the original paper. Offline methods assume access to the entire sequence and do not support local updates. The numbers in brackets indicate the TSDF Fusion resolution. While we present results for both 2cm and 4cm resolutions (with 4cm commonly used in volumetric methods), our point-based representation is not constrained by any specific TSDF Fusion resolution.
[…]
Offline COLMAP 0.069 0.135 0.102 0.634 0.505 0.558
Atlas 0.084 0.102 0.093 0.598 0.565 0.578
VoRTX 0.082 0.062 0.072 0.688 0.607 0.644
CVRecon 0.077 0.045 0.061 0.753 0.639 0.690
Online DeepVMVS 0.076 0.117 0.097 0.451 0.558 0.496
NeuralRecon 0.128 0.054 0.091 0.684 0.479 0.562
TF 0.099 0.078 0.089 0.648 0.547 0.591
SimpleRecon 0.078 0.065 0.072 0.641 0.581 0.608
VisFusion 0.105 0.055 0.080 0.695 0.527 0.598
PointRecon (4cm) 0.082 0.086 0.084 0.568 0.556 0.560
PointRecon (2cm) 0.073 0.083 0.078 0.586 0.612 0.599
Table 2: Mesh Evaluation. Mesh reconstruction quality for the ScanNetv2 test split, following the evaluation protocol of Atlas [23]. The number in parentheses represents the TSDF Fusion resolution.
@@ 4.1 Implementation Details
We select AFF-Mini, the smallest variant from [40] consisting of only 6.75M parameters, as our image encoder backbone. Our model is trained and evaluated on the ScanNetv2 [7] dataset. We resize the input resolution to $640\times 480$. The finest-level depth maps are predicted at a resolution of $160\times 120$, creating 19,200 points each new frame.
@@
Training details. We begin with training the model on image pairs for 2 epochs, using the two images to mutually serve as sources for feature matching, without any point merging at this stage. Then, we train the model on 9-view local image windows for 6 epochs, feeding the images sequentially into the model and performing point merging at each time step. We use a learning rate $10^{-4}$ for the first $4$ epochs, reducing it by a factor of 0.1 at the fifth epoch and the eighth epoch. We train with the AdamW optimizer [19] with a weight decay of $10^{-4}$.
During training, we apply random color augmentation to the input RGB images with a probability of 0.8, using TorchVision [21]. The color augmentation parameters include brightness set to 0.4, contrast to 0.4, saturation to 0.2, and hue to 0.1. Additionally, we randomly reverse the order of the input sequences with a probability of 0.5. For keyframe selection, we follow the methodology outlined in [8]. We reduce the weight for the loss of monocular depth prediction by a factor of 0.5 at the fifth and eighth epochs.
Evaluation details. During evaluation, we sequentially feed the keyframe images into the model, performing point merging at each time step. Once the point cloud for the entire scene is constructed, we render depth maps using the keyframe cameras by projecting the points onto the image planes. Each pixel in the rendered depth maps corresponds to the depth of the closest point (as detailed in Sec. 3.4). To construct a 3D mesh from the rendered depth maps, we apply TSDF fusion [6].
(a) NeuroRecon
(b) SimpleRecon
(c) Ours
(d) Ground truth
(e) NeuroRecon
(f) SimpleRecon
(g) Ours
(h) Ground truth
Figure 4: Visualization of generated meshes (please magnify when viewing). We first render depth maps from the scene point cloud and then fuse them using TSDF fusion to generate mesh. Our method produces more detailed reconstructions compared to previous work, though less smooth surfaces may result from the inherent discrete nature of point clouds. Check out our project page https://arthurhero.github.io/projects/pointrecon/ for incremental reconstruction videos.